A long horizontal wire carries 23.0 A of current due north. The Earth's field 20.0 cm due west of the wire points north but downward, 44° below the horizontal, and has magnitude of 5.0 × 10-5T


Part A

What is the magnitude of the net magnetic field 20.0 cm due west of the wire?

Part B

What is the angle below the horizontal of the net magnetic field 20.0 cm due west of the wire?
Answer #1

A)

Magnetic field, B = u0 I/2 pi r

B = (4 pi x 10^-7 x 23)/(2 pi x 0.2) = 0.000023 T

Bx = (5 x 10^-5) cos(44) = 3.6 x 10^-5 T

By = (0.000023) - (5 x 10^-5) sin(44) = - 1.17 x 10^-5 T

Magnitude of field, B' = sqrt(Bx^2 + By^2) = 3.78 x 10^-5 T

B)

Theta = arctan (By/Bx) = 18° below horizontal

Comment in case any doubt please rate my answer....